calculate the speed of sound in air at stp. density of air at STP = 1.293 kg/m^3, y = 1.40


1
Expert's answer
2021-12-07T09:54:16-0500
"v=\sqrt{\frac{1.4\cdot101325}{1.293}}=331.2\frac{m}{s}"